3. La Besurta

’TUCA DE MULLERES’ (MULLERES PEAK) 4 hours and 30 min. 3,010 m. high

Take the track at the end of the asphalt trail of the Rencllusa and Llanos del Hospital. Follow the path to the Forau d’Aiguallut, which you will reach in 30 minutes. Once there, cross and follow the Escaleta ravine, which will take you to the Pllan d’Aiguallut by its left side. After 1 hour and a half, the valley widens, descends and then narrows again. You will find the Ibons de la Escaleta (little mountain lakes) one after the other as you ascend. When you reach the last one of them, under the Facanada peak, the valley twists to the right (SW). Depending on the amount of snow, circumvent the ravine by one side or the other. You will get to another ibón under the Tuca de Barrancs, with two ravines separated by the Salenques mount. Take the left one, going uphill by its left side and you will reach the Tuca de Mulleres (3,010 m. high) after approximately 4 hours and a half. Another way of getting there is going up to the hill (2,880 m.) and from there going up by its left side to the Tuca (or peak).

High Mountain Excursion. It is essential that you be accompanied by a High Mountain Guide (if you do not have previous experience), wear mountain boots and material such as crampons and ice axe. Not recommended for children.